# The Value of AI Advice: Personalized and Value-Maximizing AI Advisors Are Necessary to Reliably Benefit Experts and Organizations

### v1

@@ A.2.1 Step 1: Simulating Human Decisions
Difficulty-Biased behavior is simulated by first training an out-of-the-box probabilistic model, such as logistic regression, to predict the correct label for unseen task instances. We can denote the probabilistic output of the model as $\hat{p}_{logistic}(y|x)$, and we use this output as a proxy for the difficulty of the instance. Specifically, we denote the proxy value for instance $i$’s difficulty with $d_{i}$, such that $d_{i}=2|\hat{p}_{logistic}(y=1|x)-0.5|$. If $d_{i}>d_{j}$, then we consider instance $i$ as easier to classify than instance $j$, because the probabilistic model predicts its class with a higher likelihood (is more confident of the instance’s label). Finally, we partition the instance space such that each value $d_{i}$ for all instances $i$ maps to the simulated human’s likelihood of being correct, and we create regions such that the human has a high accuracy for some difficult instances, and a low accuracy for some easier instances. Specifically, we set the human’s accuracy to $60\%$ whenever $d>d_{t}$ and $100\%$ whenever $d<=d_{t}$, where $d_{t}$ is a dataset specific difficulty threshold. $d_{t}$ is set to $0.6$ for the Heart Disease dataset, $0.3$ for the FICO dataset, and $0.8$ for the HR dataset. Values were chosen so that the human would have $100\%$ accuracy on a smaller part of the dataset (roughly 25%-30% of instances) and 60% on the rest, leading each human to have an overall accuracy ranging from 70%-80%.
@@
  • Heart Disease: Age $<50$
  • FICO: Number of Satisfactory Trades $<24$
  • HR: Age $>32$
@@ A.2.2 Step 2: Simulating Human Confidence in Their Decisions
  • FICO: Confidence is set to $0.2$ when external risk estimate $<65$, otherwise confidence set to $1$.
@@ A.2.3 Step 3: Simulating and Modeling Human ADB
The parameter $\gamma$ varies from 0 to 1 and controls the degree to which the human adjusts their own confidence values and the AI’s reported confidence values to $0.5$ before integrating them into a single value that represents the human’s confidence in the AI’s decision. When $\gamma=0$, all the probabilities are transformed to 0.5, while when $\gamma=1$, no adjustment is applied to probabilities before the human integrates the confidences. We select $\gamma=0.95$ to represent a human that slightly adjusts probabilities closer to 0.5. Once the human has an estimate of the AI’s probability of being correct, this probability is further transformed into a perceived probability, motivated by human tendency to perceive probabilities non-linearly [27]. The $k$ parameter determines how the likelihood they believe the AI to be correct is transformed into a perceived probability. We select a value of $k=0.63$, which corresponds to an inverse-S shaped probability weighting function, i.e. the human overweights extreme events, and underweights common events. We select this value because expert DM’s tend to overestimate the likelihood of rare diseases (rare event) and thus underestimate the lack of rare diseases (common event) [64]. Finally, given the human’s final perceived probability of the AI being correct, $w(p)$, the $\beta_{ADB}$ parameter determines how much utility they expect to gain from accepting/rejecting AI advice. The $\beta_{ADB}$ parameter can be used to increase the disparity between utility gained from decisions expected to be correct and incorrect. A $\beta_{ADB}$ value of 0 means the expected utility gained from accepting advice is equal to $w(p)$ and the expected utility gained from rejecting advice is equal to $1-w(p)$. As $\beta_{ADB}$ increases, the gap between expected utility gains increases, corresponding to higher stakes settings. We thus set $\beta_{ADB}$ to 0.5, slightly increasing the perceived stakes across all settings in our main results.
@@ A.3 Training Parameters
A.4 Measures of Performance.
@@
Appendix B Alternative Rule-Production Methods
Task-Only
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbbm{C}(R,x_{i})\\ 0,&\text{otherwise}\end{cases}$ (13)
And it offers model confidence $c^{M}_{i}$:
$c^{M}_{i}$ =$=\begin{cases}\underset{r\in R:\mathbbm{C}(r,x_{i})=1}{\max}\Bigg{[}\frac{\sum_{j=1}^{n}\big{(}\mathbbm{1}\{\mathbbm{C}(r,x_{j})\wedge\mathbbm{1}(y_{j}=\hat{y}_{i})\}\big{)}}{\sum_{k=1}^{n}\mathbbm{C}(r,x_{k})}\Bigg{]},&\text{ if }\hat{y}_{i}=1\\ \frac{\sum_{j=1}^{n}\mathbbm{1}\{y_{j}=\hat{y}_{i}\wedge y_{j}=0\}}{\sum_{k=1}^{n}\mathbbm{1}(\hat{y}_{i}=0)},&\text{ otherwise }\end{cases}$ (14)
@@ Appendix C Additional Results
In this section, we provide results on additional settings and offer additional analysis on the benefits offered by tr’s mechanisms. Specifically, we demonstrate how tr can add more value than alternatives when advising costs are especially inhibitive ($\alpha>0.5$) and when the decision loss incurred from false positives and false negatives is unequal. We provide an additional case study to assess how tr adds value relative to alternatives when an expert decision-maker has confidence that is calibrated to their decision-making strengths and weaknesses. We also assess the value of personalizing the rules of a rule-based advisor - as is done in tr’s optimization scheme - by comparing to a version of tr in which the rules are optimized only for generalization accuracy on the task, but the selective advising mechanism is subsequently applied to each piece of advice generated by those rules. Finally, we also demonstrate how the value-added by tr decreases as discretion model accuracy decreases, and we demonstrate how a simple robustness mechanism can ensure that advisors optimized using low quality discretion models will not diminish the value offered by the human alone.
@@ C.1 Prohibitive Advising Costs
C.2 Asymmetric Decision Loss
Often in high-stakes settings the loss incurred from false positives and false negatives is not equal. For example, in disease diagnostics, failing to correctly diagnose a patient with a disease is a more costly error than diagnosing that a patient has a disease when they do not. We consider an additional setting for the heart disease classification task, one in which $(\lambda_{0},\lambda_{1})=(1,3)$. This setting corresponds to false negatives contributing three times as much error as false positives, respectively. We consider an expert that exhibits a group-based decision bias; The expert has a high accuracy ($95\%$) on the young and female populations, and a low accuracy ($60\%$) on the elderly-male population. We find that tr significantly improves the human’s overall decisions more than the task-only and even the tr-no(adb) approaches in this setting, shown by the lower TTL from the Total population in Table 4. Note that all methods are optimized to account for the asymmetric decision loss. TR leads to the greatest reduction in false negatives while also leading to the lowest TTL. This result demonstrates yet another practical context in which ReV-AI advisors can add superior value over alternatives.
Table 4: Advising outcomes in terms empirical Total Team Loss (TTL) $\pm$ SE after 20 repetitions on the heart disease diagnosis task in the presence of asymmetric costs from false negatives (cost $3$) and positives (cost $1$) and when there are no reconciliation costs incurred. We pair the tr, tr-no(adb), and Task-Only advisor with a simulated individual DM. We include the tr-no(ADB) advisor to demonstrate value of leveraging ADB in high-risk, asymmetric cost settings, even when there are no reconciliation costs incurred by the DM.
TR TR-No(ADB) Task-Only Human-Alone
No Disease 0.116 $\pm$ 0.005 0.099 $\pm$ 0.003 0.150 $\pm$ 0.005 0.099 $\pm$ 0.003
Has Disease 0.115 $\pm$ 0.005 0.145 $\pm$ 0.008 0.156 $\pm$ 0.012 0.213 $\pm$ 0.007
Total 0.231 $\pm$ 0.006 0.243 $\pm$ 0.005 0.307 $\pm$ 0.015 0.313 $\pm$ 0.007
C.3 Personalizing Advising Rules
In Figure 6, we compare the performance of the tr advisors to tr-SelectiveOnly advisors across contexts. The tr-SelectiveOnly advisors are not optimized to produce value-maximizing advice, instead, they are optimized equivalently to the task-only approach. However, we apply tr’s selective advising mechanism to only offer the advice from tr-SelectiveOnly if it is expected to add value given all the pillars. Therefore, the tr-SelectiveOnly advisor will only offer advice if it is expected to add value, but what advice it can offer is not personalized to add value to the human. We find that tr can add significantly more value than tr-SelectiveOnly in some settings (Figures 6b,g,i), demonstrating the potential benefit of personalizing the rules such that what advice the model can give is sufficiently accurate and convincing to add value to the given expert. Notably, in other settings, tr leads to similar value-added compared to tr-SelectiveOnly. This demonstrates that personalizing advice to add value does not diminish value across settings, and may add value in some settings.
Figure 6: Value added by AI advisors relative to human for environments with varying trade-offs between advising costs and decision benefits for Heart Disease, FICO, and HR datasets. X axis reflects contexts in which the cost from 1/x human engagements are equivalent to the loss from one incorrect decision: higher x-values reflect that fewer human engagements are permissible to achieve a given decision benefit. Results show average value-added +- SE (shaded region) over 10 repetitions.
C.4 Additional Case Study
@@
C.5 Impact of discretion model accuracy on advising performance
ReV-AI advisors are optimized to account for human ADB by utilizing a model which estimates the human’s ADB. In all main results and case studies, we simulate real-world procedures by first learning a discretion model that estimates the human’s ADB. These models are always imperfect estimates of the humans’ true ADB. While imperfect, the models used for the main results were decent approximations of the human’s true ADB (AUC $>0.9$). In this section, we further analyze the value-added from tr’s advice as the human’s ADB deviates further from the discretion model used to optimize tr. It is critical to understand AI-advising performance under imperfect discretion estimates because, in practice, it may not be possible to learn an accurate discretion model. Further, it is critical to have mechanisms in place to ensure poor discretion models do not lead to advisors that can diminish human standalone value.
We conduct an experiment on the Heart Disease dataset in which we pair a trained tr advisor with different simulated humans. The first human exhibits the same ADB that was used to train tr, and thus the AUC of the discretion model evaluated on the human’s true ADB is $>0.9$. For the subsequent humans we add random noise to their ADB such that the discretion model used to train tr is less and less representative of the human’s true ADB. We include in the legend the AUC of the discretion model on each human’s ADB. Lower AUC values mean the human exhibits a true ADB that deviates more from what tr was trained on.
Results are shown in Figure 9. The solid lines show model performance after a robustness mechanism using validation data was applied, while the corresponding dashed lines of the same color demonstrate the result when the robustness mechanism is not applied. [^7] Note that in most cases, the dashed and solid line overlap, meaning the robustness mechanism did not activate to produce different results The robustness mechanism involves evaluating the advisor performance when paired with the human on a validation set of data, and, if the advisor is found to diminish the human’s value, the advisor is not used. In practice, assessing how the advisor performs with the individual expert on a set of low-stakes validation tasks is critical for producing reliable, value-adding advisors.

### v2

@@ A.2.1 Step 1: Simulating Human Decisions
Difficulty-Biased behavior is simulated by first training an out-of-the-box probabilistic model, such as logistic regression, to predict the correct label for unseen task instances. We can denote the probabilistic output of the model as $\hat{p}_{logistic}(y|x)$, and we use this output as a proxy for the difficulty of the instance for an AI model. AI tends to under-perform relative to humans on out-of-distribution instances, and out-performs humans on in-distribution instances [73, 74], and the logistic regression output can be used to assess the extent to which an instance is in- or out-of-distribution [68]. We thus refer to instances that are more likely to be out-of-distribution as more “difficult" for an AI. Specifically, we denote the proxy value for instance $i$’s difficulty with $d_{i}$, such that $d_{i}=2|\hat{p}_{logistic}(y=1|x)-0.5|$. If $d_{i}>d_{j}$, then we consider instance $i$ as easier to classify than instance $j$, because the probabilistic model predicts its class with a higher likelihood (is more confident of the instance’s label). Finally, we partition the instance space such that each value $d_{i}$ for all instances $i$ maps to the simulated human’s likelihood of being correct, and we create regions such that the human has a high accuracy for some difficult instances, and a low accuracy for some easier instances. Specifically, we set the human’s accuracy to $60\%$ whenever $d>d_{t}$ and $100\%$ whenever $d<=d_{t}$, where $d_{t}$ is a dataset specific difficulty threshold. $d_{t}$ is set to $0.6$ for the Heart Disease dataset, $0.3$ for the FICO dataset, and $0.8$ for the HR dataset. Values were chosen so that the human would have $100\%$ accuracy on a smaller part of the dataset (roughly 25%-30% of instances) and 60% on the rest, leading each human to have an overall accuracy ranging from 70%-80%.
@@
  • Heart Disease: Age <50<50
  • FICO: Number of Satisfactory Trades <24<24
  • HR: Age >32>32
@@ A.2.2 Step 2: Simulating Human Confidence in Their Decisions
  • FICO: Confidence is set to $0.2$ when external risk estimate <65<65, otherwise confidence set to $1$.
@@ A.2.3 Step 3: Simulating and Modeling Human ADB
The parameter $\gamma$ varies from 0 to 1 and controls the degree to which the human adjusts their own confidence values and the AI’s reported confidence values to $0.5$ before integrating them into a single value that represents the human’s confidence in the AI’s decision. When $\gamma=0$, all the probabilities are transformed to 0.5, while when $\gamma=1$, no adjustment is applied to probabilities before the human integrates the confidences. We select $\gamma=0.95$ to represent a human that slightly adjusts probabilities closer to 0.5. Once the human has an estimate of the AI’s probability of being correct, this probability is further transformed into a perceived probability, motivated by human tendency to perceive probabilities non-linearly [33]. The $k$ parameter determines how the likelihood they believe the AI to be correct is transformed into a perceived probability. We select a value of $k=0.63$, which corresponds to an inverse-S shaped probability weighting function, i.e. the human overweights extreme events, and underweights common events. We select this value because expert DM’s tend to overestimate the likelihood of rare diseases (rare event) and thus underestimate the lack of rare diseases (common event) [85]. Finally, given the human’s final perceived probability of the AI being correct, $w(p)$, the $\beta_{ADB}$ parameter determines how much utility they expect to gain from accepting/rejecting AI advice. The $\beta_{ADB}$ parameter can be used to increase the disparity between utility gained from decisions expected to be correct and incorrect. A $\beta_{ADB}$ value of 0 means the expected utility gained from accepting advice is equal to $w(p)$ and the expected utility gained from rejecting advice is equal to 1w(p)1-w(p). As $\beta_{ADB}$ increases, the gap between expected utility gains increases, corresponding to higher stakes settings. We thus set $\beta_{ADB}$ to 0.5, slightly increasing the perceived stakes across all settings in our main results.
@@ A.3 Training Parameters
Appendix B Alternative Rule-Production Methods
B.1 Task-Only
$\hat{y}_{i}=\begin{cases}1,&\text{if }\mathbb{C}(R,x_{i})\\ 0,&\text{otherwise}\end{cases}$ (13)
And it offers model confidence $c^{M}_{i}$:
@@ Appendix C Additional Results
In this section, we provide results on additional settings and offer additional analysis on the benefits offered by TR’s mechanisms. Specifically, we demonstrate how TR can add more value than alternatives when advising costs are especially inhibitive ($\alpha>0.5$) and when the decision loss incurred from false positives and false negatives is unequal. We provide an additional case study to assess how TR adds value relative to alternatives when an expert decision-maker has confidence that is calibrated to their decision-making strengths and weaknesses. Finally, we also demonstrate how the value-added by TR decreases as discretion model accuracy decreases, and we demonstrate how a simple robustness mechanism can ensure that advisors optimized using low quality discretion models will not diminish the value offered by the human alone.
@@ C.1 Prohibitive Advising Costs
C.2 Additional Case Study
@@
C.3 Impact of discretion model accuracy on advising performance
In this section, we introduce and assess the benefit of a robustness mechanism that aims to further safeguard an AI advisor from offering advice that might diminish the human’s own value. Concretely, the simple mechanism involves an empirical evaluation of the AI advisor’s value implications when paired with the human on a validation set; if the advisor is found empirically to diminish the human’s value, the advisor is not used. In practice, assessing how the advisor performs with the individual expert on a set of risk-free validation tasks is instrumental for producing reliable, value-adding advisors.
ReV-AI advisors are designed to complement the human ADB by utilizing a discretion model which estimates the human’s ADB. In all main results and case studies, we simulate real-world procedures by first learning a discretion model from human-AI interaction data, which produces imperfect estimates of the human’s true ADB. While imperfect, the discretion models produced in the main results were decent approximations of the human’s true ADB (AUC >0.9>0.9). In this section, we further analyze the value-added from TR’s advice as the human’s ADB deviates further from the discretion model used to optimize TR. It is critical to understand AI-advising performance under imperfect discretion estimates because, in practice, it may not be possible to learn an accurate discretion model. Further, it is critical to have mechanisms in place to ensure poor discretion models do not lead to advisors that can diminish human standalone value.
We conduct an experiment on the Heart Disease dataset in which we pair a trained TR advisor with different simulated humans. The first human exhibits the same ADB that was used to train TR, and thus the AUC of the discretion model evaluated on the human’s true ADB is >0.9>0.9. For the subsequent humans we add random noise to their ADB such that the discretion model used to train TR is less and less representative of the human’s true ADB. We include in the legend the AUC of the discretion model on each human’s ADB. Lower AUC values mean the human exhibits a true ADB that deviates more from what TR was trained on.
Results are shown in Figure 10. The solid lines show model performance after a robustness mechanism using validation data was applied, while the corresponding dashed lines of the same color demonstrate the result when the robustness mechanism is not applied. [^8] Note that in most cases, the dashed and solid line overlap, meaning the robustness mechanism did not activate to produce different results